OpenStudy (anonymous):

HELP! pleaseeee!!!! I WILL GIVE MEDAL! AB and BC form a right angle at point B. If A =(-3,-1) and B =(4, 4), what is the equation of BC?

OpenStudy (anonymous):

A. x + 3y = 16 B. 2x + y = 12 C. -7x − 5y = -48 D. 7x − 5y = 48

OpenStudy (anonymous):

I need help here this is completely confusing me... :/

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

all so far i did is graph but otherwise now im lost

OpenStudy (anonymous):

what grade r u in

OpenStudy (anonymous):

Im in 12th grade.

OpenStudy (anonymous):

ok then. sry i'm only in 6th. I don't know this

OpenStudy (anonymous):

@iGreen @AnswerMyQuestions @thomaster @aaronq Please any of you guys if you could help me and lead me. I don't want a direct answer I want to understand this that way I don't have the same problem with similar questions.

OpenStudy (anonymous):

I have no clue where C is located..

OpenStudy (anonymous):

@iGreen @perl can you help?

OpenStudy (anonymous):

@ajprincess @King.Void.

OpenStudy (anonymous):

I just need an understanding of this not a direct answer

OpenStudy (anonymous):

I know and @iGreen and @perl can help with that :)

OpenStudy (anonymous):

@MSxSammiBooxX could probably help to

OpenStudy (anonymous):

thanks for coming @iGreen

OpenStudy (anonymous):

@owllover123 thank you for helping out :)

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

I need help now I have to finish this online test before the bell rings, I already got a similar question wrong and I don't want that. I need help

OpenStudy (anonymous):

@iGreen can you help?

OpenStudy (anonymous):

Im leaving in 10 minutes exact. Im just going to have to guess it by 11:16 to finish the other questions

OpenStudy (anonymous):

Do you have a graph or image??

OpenStudy (anonymous):

It doesnt include that thats why Im confused

OpenStudy (anonymous):

Ok, well we are going to graph it ourselves. Give me a few seconds to do so

OpenStudy (anonymous):

I graphed it on my paper to visualize it

OpenStudy (anonymous):

i got 6mins left :/

OpenStudy (anonymous):

Well, we have these points http://prntscr.com/6rm9tk for it to be a right triangle, the other point must be located in one of these two spots

OpenStudy (anonymous):

Hold on let me label the coordinates

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

http://prntscr.com/6rmb4a Ok, let's look at the C point located above the line I drew

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

If we look at that, BC would be a straight, horizontal line, so what would the slope of that be?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

I have a minute left :(

OpenStudy (anonymous):

please keep explaining. If i dont pass the test I can still review this for help!@ but thaqnk you so much

OpenStudy (anonymous):

I will resume tomorrow

OpenStudy (anonymous):

Ah sorry I gotta go, @iGreen can you continue? I am getting my hair colored right now and my timer thing went off, so I gotta go and get this dye out

OpenStudy (ajprincess):

it is given that AB and BC meet at right angles. that means they are perpendicular to each other. Then the products of the slopes of the two lines will be equal to -1. Let the slope of AB be \(m_1\) and the slope of BC be \(m_2\). \(m_1*m_2=-1\). \(m_1\) can be found using the slope formula. It is \[m_1=\frac{ y_2-y_1 }{ x_2-x_1}\] where \(m_1\)-slope \((x_1,y_1)\) and \((x_2,y_2)\) - coordinates of the two points

OpenStudy (ajprincess):

\(m_2\) can be found by substituting the value of \(m_1\) that was found from the slope formula in \(m_1*m_2=-1\)

OpenStudy (ajprincess):

Let the equation of BC be y=mx+c m-slope of the line. c-intercept. m=\(m_2\) Substituting the value of \(m_2\) and \(x_2,y_2\) in the above equation c can be found. i.e \(y_2=m_2x_2+c\) . Here, \((x_2,y_2)\) is the coordinate of B. Then substituting the values of m_2 and c in the equation y=mx+c will give you the required equation.

OpenStudy (ajprincess):

If you need further explanation pls feel free to ask me.:)